OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. x - y = 120 x + y = 150

OpenStudy (anonymous):

Addition is the easier way to solve for x here, but substitution would work as well. For addition, since we know the values for x and y that satisfy these equations work for both, we can add the two equations and keep the equality. So, take the vertical sum: x+x-y+y=120+150, which simplifies to: 2x=270. Dividing by 2: x=135. Incidentally, y =15.